OpenStudy (anonymous):

√27x^3y^4 OVER √3xy^2

OpenStudy (anonymous):

Like so?\[\frac{\sqrt{27}x^3y^4}{\sqrt{3}xy^2}\]

OpenStudy (anonymous):

@jabberwock yessirr!

OpenStudy (anonymous):

Okidoki. So let's separate this a little bit to make what we're going to do easier to see: \[\frac{\sqrt{27}}{\sqrt{3}}\frac{x^3}{x}\frac{y^4}{y^2}\] If we only have multiplication or division of square roots, we can combine the square roots like this: \[\sqrt{\frac{27}{3}}\frac{x^3}{x}\frac{y^4}{y^2}\]

OpenStudy (anonymous):

The numerical part simplifies to the square root of 9, which is 3 \[3\frac{x^3}{x}\frac{y^4}{y^2}\]Good so far?

OpenStudy (unklerhaukus):

\[\frac{a^n}{a^m}=a^{n-m}\]

OpenStudy (anonymous):

yesss soooo sorry @jabberwock continue!

OpenStudy (anonymous):

Alright. Now consider just the part with the x's: \[\frac{x^3}{x}=\frac{x*x*x}{x}=x^2\]

OpenStudy (anonymous):

why did that just turn into x squared?

OpenStudy (anonymous):

Ah, one of the x's up top cancels with one of the x's on the bottom

OpenStudy (anonymous):

So you just have x*x after that

OpenStudy (anonymous):

ohhhh that makes sense! haha okay what next?

OpenStudy (anonymous):

Same reasoning with the y's \[\frac{y^4}{y^2}=\frac{y*y*y*y}{y*y}=y^2\]

OpenStudy (anonymous):

We are cancelling two y's up top with two y's on the bottom

OpenStudy (anonymous):

okayy cool. so we have y squared and x squared.. so the final would be 3x^2y^2???

OpenStudy (anonymous):

So, quick note: \[x=x^1\] Look at the x's again: \[\frac{x^3}{x^1}=x^{3-1}=x^2\]Look at the y's: \[\frac{y^4}{y^2}=y^{4-2}=y^2\]

OpenStudy (anonymous):

wow okay, just simple subtraction!

OpenStudy (anonymous):

So as a shortcut, you can just subtract when you have exponents with the same base on the top and the bottom. For example: \[\frac{x^{100}}{x^{75}}=x^{25}\] And yes, that final answer you gave looks right :)

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

Sorry, looks like I'm lagging

OpenStudy (anonymous):

haha its fine! thanks soo much! i have more tho!

OpenStudy (anonymous):

last one!

OpenStudy (anonymous):

No worries. K, shoot.

OpenStudy (anonymous):

\[3\sqrt{24}+\sqrt{54}-\sqrt{6}\]

OpenStudy (anonymous):

Okidoki. This is slightly more complicated. Let's just deal with the first term. \[3\sqrt{24}\]Let's look for a PERFECT SQUARE that divides into 24

OpenStudy (anonymous):

Ideas?

OpenStudy (anonymous):

ohhhh! √6 x √4

OpenStudy (anonymous):

Exactly! Know what to do with it?

OpenStudy (anonymous):

well do i simplify the √6 even more? and the other would be 2..

OpenStudy (anonymous):

2 isn't a perfect square, so we can't use that.

OpenStudy (anonymous):

We're only going to pull out perfect squares. If we could have pulled out a 4 or a 9, we would have done that, too.

OpenStudy (anonymous):

\[3\sqrt{6}(\sqrt{4})=3\sqrt{6}(2)=6\sqrt{6}\]

OpenStudy (anonymous):

Okay with that?

OpenStudy (anonymous):

i dont understand that..

OpenStudy (anonymous):

ohhhh nevermind i get it

OpenStudy (anonymous):

Cool cool. Want to try the next term?

OpenStudy (anonymous):

uhmm could it be √6+√9?

OpenStudy (anonymous):

so it'd be √6+3?

OpenStudy (anonymous):

\[\sqrt{6}\sqrt{9}=3\sqrt{6}\]Multiplication instead of addition.

OpenStudy (anonymous):

alright so is that as far as we can simplify?

OpenStudy (anonymous):

We now have \[6\sqrt{6}+3\sqrt{6}-\sqrt{6}\] Think of the square root of 6 like a variable. You can combine them kind of like like-terms.

OpenStudy (anonymous):

Let me know if you need help with that part.

OpenStudy (anonymous):

so we divide both sides by 3?

OpenStudy (anonymous):

There's only one side, because this isn't an equation (there's no equal sign)

OpenStudy (anonymous):

How would you simplify 6x+3x-x?

OpenStudy (anonymous):

8x.

OpenStudy (anonymous):

so it would be 8√6!

OpenStudy (anonymous):

Yep. So simplify \[6\sqrt6+3\sqrt{6}-\sqrt{6}\]

OpenStudy (anonymous):

Exactly :)

OpenStudy (anonymous):

so what would it be when we put it all together?

OpenStudy (anonymous):

That's it. 8root6

OpenStudy (anonymous):

okay thanks so much!

OpenStudy (anonymous):

No problem!